Calculate the mass of 1.505 molecules of oxygen gas
1
Expert's answer
2019-09-30T04:52:56-0400

First, the number of moles of oxygen gas should be estimated:

n (O2) = N(O2) / NA,

where n - number of moles, N - number of oxygen gas molecules, NA - Avogadro's number.

From here:

n (O2) = 1505 / 6.022 × 1023 = 249.92 × 10-23 moles

The mass of the oxygen gas is:

m (O2) = M (O2) × n (O2),

where m - mass, M - molecular weight, n - number of moles.

From here:

m (O2) = 32 g/mol × 249.92 × 10-23 moles = 8 × 10-20 g.


Answer: 8 × 10-20 g